A person sitting on the top of a tall building is dropping balls at regular intervals of one second. Find the positions of the 3rd. 4th and 5th ball when the 6th ball is being dropped.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. For every ball, u = 0, a = g = 9.8 m/s2 ∴ 4th ball move for 2 sec, 5th ball 1 sec and 3rd ball 3 sec when 6th ball is being dropped. For 3rd ball t = 3 sec S3 = ut + ½ at2 0 + ½ (9.8) 32 = 4.9 m below the top. For 4th ball, t = 2 sec S2 = 0 + 1/2 gt2 = 1/2 (9.8)22 = 19.6 m below the top (u = 0) For 5th ball, t = 1 sec S3 = ut + 1/2 at2 = 0 + 1/2 (9.8)t2 = 4.98 m below the top.